OpenStudy (anonymous):

how do u mult fractions??? i typed in 15/16 x 4/9 and it came out as 5/1

OpenStudy (anonymous):

5/12 sorry

OpenStudy (anonymous):

when multiplying fractions, you can simply multiply across the two, so: 15/16 * 4/9 = (15 * 4)/(16 * 9) = 60/144 We can further reduce this new fraction by pulling a 12 out of the top and bottom: 12(5)/12(12), cancelling the 12, we are left with 5/12.

OpenStudy (lgbasallote):

the product of the numerators over the product of the denominators

OpenStudy (anonymous):

where do you get 12 from both top and bottom?

OpenStudy (anonymous):

never mind.